OpenStudy (anonymous):

Given a = 4, b = -2 and c = 8, evaluate ab - bc. is it -4?

OpenStudy (anonymous):

-8+16=8

OpenStudy (anonymous):

4*(-2)=-8 (-2)*8=-16 Sooo -8+16=8